# LLM-based User Profile Management for Recommender System

### v1

@@ Abstract
The rapid advancement of Large Language Models (LLMs) has opened new opportunities in recommender systems by enabling zero-shot recommendation without conventional training. Despite their potential, most existing works rely solely on users’ purchase histories, leaving significant room for improvement by incorporating user-generated textual data, such as reviews and product descriptions. Addressing this gap, we propose PURE, a novel LLM-based recommendation framework that builds and maintains evolving user profiles by systematically extracting and summarizing key information from user reviews. PURE consists of three core components: a Review Extractor for identifying user preferences and key product features, a Profile Updater for refining and updating user profiles, and a Recommender for generating personalized recommendations using the most current profile. To evaluate PURE, we introduce a continuous sequential recommendation task that reflects real-world scenarios by adding reviews over time and updating predictions incrementally. Our experimental results on Amazon datasets demonstrate that PURE outperforms existing LLM-based methods, effectively leveraging long-term user information while managing token limitations.
LLM-based User Profile Management for Recommender System
Seunghwan Bang1     Hwanjun Song2,† 1Ulsan National Institute of Science and Technology 2Korea Advanced Institute of Science and Technology shbang1422@unist.ac.kr songhwanjun@kaist.ac.kr
$\dagger$ indicates the corresponding author.
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) Touvron et al. (2023); Dubey et al. (2024); Achiam et al. (2023); Team et al. (2024) has significantly impacted various domains, such as text summarization Lewis et al. (2020a) and search Karpukhin et al. (2020). Recent studies leverage LLMs in recommender systems for their human-like reasoning and external knowledge integration through in-context learning Brown et al. (2020) and retrieval-augmented generation Lewis et al. (2020b). As such, LLMs exhibit the potential to be used as zero-shot recommendation models without conventional training, which traditionally relies on explicit user-item interactions and training data He et al. (2017); Kang and McAuley (2018); He et al. (2020).
Figure 1: Overall system of PURE. PURE incorporates reviews, ratings, and item interactions, whereas LLM Recommender handles only item interactions. By using the "Review Extractor" to identify key information and the "Profile Updater" to refine the user profile, PURE addresses scalability issue (i.e., growth of input token size).
@@
Despite the advanced capability of LLMs, most recent works Hou et al. (2024); Wei et al. (2024); Ren et al. (2024); He et al. (2023); Zhai et al. (2023) rely solely on users’ past purchase history (i.e., list of purchased items). This leaves significant room for further improvement by incorporating additional user-generated textual information, such as user reviews and product descriptions, which have yet to be fully leveraged. In other words, they still fail to fully leverage various text data due to their inability to retain and process the increasing contextual information as users continue to make purchases, leading to longer recommendation sessions. This issue is primarily attributed to the omission of the context, either due to the information loss within the LLM’s memory Liu et al. (2024) or the memory capacity by the token limit Li et al. (2024); Ding et al. (2024). Thus, extracting key features from a user’s diverse textual sources is essential, as demonstrated in MemoryBank Zhong et al. (2024), a framework that enhances LLMs with long-term memory by summarizing key information from conversations and updating user profiles.
@@
Building on this foundation, we take the first step in extending LLMs’ long-term memory beyond conversations in MemoryBank, adapting it to the evolving dynamics of recommendation systems. We propose PURE, a novel LLM-based Profile Update for REcommender that constructs user profile by integrating users’ purchase history and reviews, which naturally expand as the recommendation sessions progress. Designed specifically for recommendation in Fig. 1, PURE systematically extracts user preferences, dislikes, and key features from reviews and integrates them into structured user profiles. Specifically, PURE consists of three main components: "Review Extractor", which analyzes user reviews to identify and extract user preferences, dislikes, and preferred product features, referred to as "key features", offering a comprehensive view of user interests and purchase-driving attributes; "Profile Updater", which refines newly extracted representations by eliminating redundancies and resolves conflicts with the existing user profile, ensuring a compact and coherent user profile; and "Recommender", which utilizes the most up-to-date user profile for recommendation task.
@@
Our main contributions are as follows: (1) We propose PURE, a novel framework that systematically extracts, summarizes, and stores key information from user reviews, optimizing LLM memory management for the recommendation. (2) We validate the effectiveness of PURE by introducing a more realistic sequential recommendation setting, where reviews are incrementally added over time, allowing the model to update user profiles and predict the next purchase continuously. This setup more accurately reflects real-world recommendation scenarios compared to prior works, which assume all past purchases are provided at once, ignoring the evolving nature of user preferences. (3) We empirically show that PURE surpasses existing LLM-based recommendation methods on Amazon data, demonstrating its effectiveness in leveraging lengthy purchase history and user reviews.
2 Related Works
Recommendation Setup.
Conventional sequential recommendation methods Wang et al. (2019); Kang and McAuley (2018); Sun et al. (2019); Hidasi and Karatzoglou (2018); Kim et al. (2024) followed a one-shot prediction setup, where user history is split: the last item as the test set, the second-to-last as validation, and the rest for training. These models predict a single target item, failing to capture evolving user behavior. Tallrec Bao et al. (2023) framed the task as binary classification to predict whether an item should be recommended.
LLM-based Recommendation.
Tallrec Bao et al. (2023) proposed the parameter efficient finetuning (PEFT) method in recommendation system, and A-LLMRec Kim et al. (2024) proposed to finetune the embedding model for LLM to leverage the collaborative knowledge. In contrast, LLM elicited responses and extracted multiple representations from the conversation without extra training Wang and Lim (2023). Moreover, the authors Dai et al. (2023) showed the potential of ChatGPT for reranking the candidates. InstructRec Zhang et al. (2023) designed the instruction to recognize the users’ intention and preference from context.
@@ 3.2 PURE: Profile Update for REcommender
Input: Review extractor $\mathcal{E(\cdot)}$, User profile updater $\mathcal{U}(\cdot)$, Recommender $\mathcal{R}(\cdot)$, Dataset $\mathfrak{D}_{u}=\{\mathbf{R}_{u},\mathbf{I}_{u}\}$ for user $u$, User profile $\mathbf{P}_{u}^{t}$, next purchase candidates $\mathcal{C}_{u}^{t+1}$, timestep $t$
@@
In this section, we introduce PURE, novel framework that manages the user profile $\mathbf{P}_{u}$ from user reviews $\mathbf{R_{u}}$ and predict the next item with user profile. Algorithm 1 can be divided into three steps (See Appendix A for prompt template).
STEP 1: Extract User Representation.
@@
We begin by providing the LLM with raw inputs, including user reviews $\mathbf{R_{u}}$ and product names $\mathbf{I_{u}}$. The LLM extracts $\tilde{l_{u}^{t}}$(items the user likes), $\tilde{d_{u}^{t}}$(items the user dislikes), and $\tilde{f_{u}^{t}}$(key user features) from the incoming review as user representation.
STEP 2: Update User Profile.
@@
After the extraction in STEP 1, the extracted representation <$\tilde{l_{u}^{t}}$, $\tilde{d_{u}^{t}}$, $\tilde{f_{u}^{t}}$> concatenates with previous user profile $\mathbf{P}_{u}^{t-1}=\{l_{u}^{t-1},d_{u}^{t-1},f_{u}^{t-1}\}$. However, this faces a scalability issue as the number of reviews increases. Thus, leveraging the previous profile, we use an LLM to remove redundant and conflicting content from the extracted representation, yielding a more compact and up-to-date user profile $\mathbf{P}_{u}^{t}$ after concatenation.
STEP 3: Recommend Next Purhcase Item.
@@
Recommender $\mathcal{R}$ reranks the given candidate item list to predict the user’s next purchase by leveraging the updated profile $\mathbf{P}_{u}^{t}$ and purchased items $\mathbf{I}_{u}$.
@@
Table 1: Comparison PURE with Baselines. We evaluate performance under two data settings: using only item interactions and using item interactions augmented with reviews. $\dagger$ indicates customized baselines where review data is naively incorporated into the original prompt templates designed for item interactions only (see  Sec. C.2).
@@
4 Experiment
Datasets. For a thorough evaluation, we utilize two datasets from the Amazon collection Ni et al. (2019): Video Games and Movies & TV. To ensure a comprehensive analysis, we intentionally select datasets with diverse statistical properties, particularly in terms of the number of items (See Appendix B for details).
Baselines.  Hou et al. (2024) is the recommendation method that utilizes pre-trained LLMs without additional training or fine-tuning, making it a suitable baseline. It describes three approaches for LLM-based recommendation: Sequential, Recency, and in-context learning (ICL). We compare our method with all three approaches and demonstrate the superiority of PURE when these techniques were applied to our framework, further highlighting its effectiveness. (See Sec. C.1 for details.)
Evaluation Setting. To assess the performance of PURE, we adopt a continuous sequential recommendation task. Note that NDCG scores are first aggregated per user across multiple recommendation sessions and then across all users, reflecting the continuous nature of our setup.
Implementation Details. The prediction process is framed as a classification task where the model selects one item from candidate set $\mathcal{C}_{u}$. Each candidate set consists of 19 randomly selected non-interacted items and a ground truth item. We adopt Llama-3.2-3B-Instruct Touvron et al. (2023) as the backbone model for all the experiments.
4.1 Experimental Results
Impact of Review Extractor. Tab. 1 compares PURE with (1) three baselines solely based on purchased items; (2) modified baselines, marked with $\dagger$, that additionally utilize users’ raw reviews. The results reveal that baselines that simply combine item interactions with raw reviews show inconsistent performance improvements. In contrast, PURE, which leverages the review extractor and profile updater, significantly outperforms all baselines. This demonstrates that processing reviews at three levels, i.e., like, dislike, and key features, is essential for enhancing performance.
@@ 4.1 Experimental Results
Figure 2: Trade-off between NDCG and token size.
Component-wise Study. Tab. 2 shows the ablation study of PURE, where we analyze the impact of reviews (using or not using) and the effect of components (enabling or disabling the review extractor and profile updater). The use of reviews bring high performance gains only when accompanied by Review Extractor (Ext.). This is due to the sharp increase in input tokens (see the |$T$| column of the 2nd and 3rd rows of each method) as the user continues purchases.
Notably, the best recommendation performance is achieved when Profile Updater (Upd.) is disabled (see the 3rd and 4th rows for each method). That is well-formed context by Review Extractor can bring higher gains when simply concatenated. However, it may face a challenge, as the number of purchases grows, leading to significant computational overhead. Thus, we use the Profile Updater (Upd.) to maintain compact user profiles, reducing input token size by 15–20% with only a slight 1–3% performance drop. This trade-off underscores the importance of using Profile Updater for long-term recommendations.
Trade-off Analysis. We categorize users into three groups based on the total cumulative review token count per user, as the criterion: 0–500 (short), 500–1000 (middle), and 1000–2000 (long) tokens. Fig. 2 presents the trade-off between recommendation performance and input token length of the three models including PURE.
PURE achieves the best trade-off, showing the steepest NDCG increase compared to other methods as input token size grows. Therefore, this demonstrates that PURE accurately distills key information from long reviews, while achieving high efficiency by minimizing input token growth without information loss, even for long-group users.
@@ 5 Conclusion
We present PURE, a novel framework for LLM-based recommendation that builds and maintains evolving user profiles by systematically extracting and summarizing user representations from reviews. By introducing a continuous sequential recommendation task, we demonstrated how updating user profiles improves recommendation quality while addressing token limitation challenges.
6 Limitations
@@
A notable limitation of our approach is the tendency of the LLM to exhibit hallucination by occasionally recommending items beyond the predefined candidate set, even when explicitly instructed to select from it. This phenomenon underscores the inherent difficulty in imposing strict constraints within LLM-based recommendation models while maintaining flexibility and accuracy. Also, our study was constrained by the inability to utilize datasets containing a larger number of user reviews, which may have provided richer context.
7 Potential Risks
A potential risk associated with our approach is the possibility that user reviews may contain personal information, making data management and privacy protection critical concerns. Ensuring secure handling and anonymization of such data is essential to prevent breaches of user privacy.
8 Ethical Statement
This study used Amazon datasets which is publicly available. The dataset does not contain any personal identifiable information (PII), ensuring user privacy and ethical compliance. To ensure fair and accurate evaluation, we customized the baseline models by incorporating both item interactions and user reviews, enabling a more balanced comparison with our proposed approach. Lastly, our research aims to advance the development of recommendation systems while avoiding potential negative impacts such as bias or misuse.
Acknowledgments

### v2

@@ Abstract
The rapid advancement of Large Language Models (LLMs) has opened new opportunities in recommender systems by enabling recommendations without conventional training. Despite their potential, many existing works rely solely on users’ purchase histories, leaving significant room for improvement by incorporating user-generated textual data, such as reviews and product descriptions. Addressing this gap, we propose PURE, a novel LLM-based recommendation framework that builds and maintains evolving user profiles by systematically extracting and summarizing key information from user reviews. PURE consists of three core components: a Review Extractor for identifying user preferences and key product features, a Profile Updater for refining and updating user profiles, and a Recommender for generating personalized recommendations using the most current profile. To evaluate PURE, we introduce a continuous sequential recommendation task that reflects real-world scenarios by adding reviews over time and updating predictions incrementally. Our experimental results on Amazon datasets demonstrate that PURE outperforms existing LLM-based methods, effectively leveraging long-term user information while managing token limitations.
keywords:
Large Language Models (LLMs), Recommender Systems (RS), Personalization
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) [1, 2, 3, 4] has significantly impacted various domains, such as text summarization [5] and search [6]. Recent studies leverage LLMs in recommender systems for their human-like reasoning and external knowledge integration through in-context learning [7] and retrieval-augmented generation [8]. As such, LLMs exhibit the potential to be used as train-free recommendation models without conventional training, which traditionally relies on explicit user-item interactions and training data [9, 10, 11].
@@
Despite the advanced capability of LLMs, most recent works [12, 13, 14, 15, 16] rely solely on users’ past purchase history (ie list of purchased items). This leaves significant room for further improvement by incorporating additional user-generated textual information, such as user reviews and product descriptions, which have yet to be fully leveraged. In other words, they still fail to fully leverage various text data due to their inability to retain and process the increasing contextual information as users continue to make purchases, leading to longer recommendation sessions. This issue is primarily attributed to the omission of the context, either due to the information loss within the LLM’s memory [17] or the memory capacity by the token limit [18, 19]. Thus, extracting key features from a users’ textual sources is essential, as demonstrated in MemoryBank [20], a framework that enhances LLMs with long-term memory by summarizing key information from conversations and updating user profiles.
Figure 1: Overall system of PURE. PURE incorporates reviews, ratings, and item interactions, whereas LLM Recommender handles only item interactions. By using the "Review Extractor" to identify key information and the "Profile Updater" to refine the user profile, PURE addresses scalability issue (ie growth of input token size).
@@
Building on this foundation, we take the first step in extending LLMs’ long-term memory beyond conversations in MemoryBank, adapting it to the evolving dynamics of recommendation systems. We propose PURE, a novel LLM-based Profile Update for REcommender that constructs a user profile by integrating users’ purchase history and user-generated reviews, which naturally expand as the recommendation sessions progress. As illustrated in  Figure 1, PURE systematically extracts user likes, dislikes, and key features from reviews and integrates them into structured, dynamic user profiles. Specifically, PURE consists of three main components: "Review Extractor", which analyzes user reviews to identify and extract user likes, dislikes, and preferred product features, referred to as "key features", offering a comprehensive view of user interests and purchase-driving attributes; "Profile Updater", which refines newly extracted representations by eliminating redundancies and resolves conflicts with the existing user profile, ensuring a compact and coherent user profile; and "Recommender", which utilizes the most up-to-date user profile for recommendation task.
@@
Our main contributions are as follows: (1) We propose PURE, a novel framework that systematically extracts and stores key information from user reviews, optimizing LLM memory management for the recommendation. (2) We validate the effectiveness of PURE by introducing a more realistic sequential recommendation setting, where reviews are incrementally added over time, allowing the model to update user profiles and predict the next purchase continuously. This setup more accurately reflects real-world recommendation scenarios compared to prior works, which assume all past purchases are provided at once, ignoring the evolving nature of user preferences. (3) We empirically show that PURE surpasses existing LLM-based recommendation methods on Amazon data, demonstrating its effectiveness in leveraging lengthy purchase history and user reviews.
@@ 2.2 PURE: Profile Update for REcommender
In this section, we introduce PURE, novel framework that manages the user profile $\mathbf{P}_{u}$ from user reviews $\mathbf{R_{u}}$ and predict the next item with user profile. Algorithm 1 can be divided into three steps.
1
Input: Review extractor $\mathcal{E}(\cdot)$, User profile updater $\mathcal{U}(\cdot)$,
            Recommender $\mathcal{R}(\cdot)$, Dataset $\mathfrak{D}_{u}=\{\mathbf{R}_{u},\mathbf{I}_{u}\}$ for user $u$,
            User profile $\mathbf{P}_{u}^{t}$, next purchase candidates $\mathcal{C}_{u}^{t+1}$, timestep $t$
@@
STEP 1: Extract User Representation. We begin by providing the LLM with raw inputs, including user reviews $\mathbf{R_{u}}$ and product names $\mathbf{I_{u}}$. The LLM extracts $\tilde{l_{u}^{t}}$(items the user likes), $\tilde{d_{u}^{t}}$(items the user dislikes), and $\tilde{f_{u}^{t}}$(key user features) from the incoming review as user representation. To do so, we utilized the following prompt template:
I purchased the following products and left reviews in chronological order: {Asins, product names, input reviews}. Analyze user’s likes/dislikes/key features by referring to their reviews.
@@
STEP 2: Update User Profile. After the extraction in STEP 1, the extracted representation <$\tilde{l_{u}^{t}}$, $\tilde{d_{u}^{t}}$, $\tilde{f_{u}^{t}}$> concatenates with previous user profile $\mathbf{P}_{u}^{t-1}=\{l_{u}^{t-1},d_{u}^{t-1},f_{u}^{t-1}\}$. However, this faces a scalability issue as the number of reviews increases. Thus, leveraging the previous profile, we use an LLM to remove redundant and conflicting content from the extracted representation, yielding a more compact and up-to-date user profile $\mathbf{P}_{u}^{t}$ after concatenation. To achieve this, we utilized the following prompt template:
You are given a list: {list}. Update this list by removing redundant or overlapping information. Note that crucial information should be preserved.
@@
STEP 3: Recommend Next Purchase Item. Recommender $\mathcal{R}$ reranks the given candidate item list to predict the user’s next purchase by leveraging the updated profile $\mathbf{P}_{u}^{t}$ and purchased items $\mathbf{I}_{u}$. As such, here is the prompt template that we utilized:
Positive aspects: {likes} Negative aspects: {dislikes} Key Features: {key features} Based on these inputs, rank the {candidate list} from 1 to 20 by evaluating their likelihood of being purchased.
3 Experiment
Datasets. For a thorough evaluation, we utilize two datasets from the Amazon collection [21]: Video Games and Movies & TV. To ensure a comprehensive analysis, we select datasets with diverse statistical properties, particularly in terms of the number of items. Each dataset includes ASINs, product names, and user reviews, which are chronologically sorted per user to reflect real-world behavior.
Baselines. LLMRank [12] is the recommendation method that utilizes pre-trained LLMs without additional training or fine-tuning, making it a suitable baseline. It describes three approaches for LLM-based recommendation: Sequential, Recency, and In-context learning (ICL). We compare our method with all three approaches and demonstrate the superiority of PURE when these techniques were applied to our framework, further highlighting its effectiveness. In the following, we describe each approach:
1) Sequential. We provide the LLM with instructions, supplying only the user-item interactions and the candidate list. The LLM is then tasked with ranking the items in the candidate list based on the likelihood of being purchased at time step $t$.
2) Recency-Focused. In the sequential prompt above, we add an instruction to emphasize the most recently purchased item, specifically the one bought at time step ($t{-}1$). The additional prompt is: "Note that my most recently purchased item is {recent item}."
3) In-Context Learning. Unlike the previous sequential and recency-focused prompts, this approach utilizes user-item interactions only up to time step ($t{-}2$) and the recently purchased item at ($t{-}1$). The additional prompt is: "I’ve purchased the following products: {user-item interactions}, then you should recommend {recent item} to me, and now that I’ve bought {recent item}."
Evaluation Setting. To evaluate the performance of PURE, we adopt a continuous sequential recommendation. In this setup, the LLM is tasked with predicting the item a user is most likely to purchase at time step $t$. The model receives the user’s interaction history up to time step $(t{-}1)$ in chronological order, along with a candidate set comprising one ground-truth item and 19 randomly sampled non-interacted items. Here, time step $t$ spans from the user’s 4th purchase to their final purchase $k$. To reflect the sequential nature of the task, NDCG scores are first aggregated across multiple recommendation sessions for each user and then averaged across all users.
Implementation Details. To perform this framework, we utilize Llama-3.2-3B-Instruct [1] as the backbone model for all experiments. Due to the inherent nature of generative language models, it is not always guaranteed that the output will follow the desired format in every response. This issue can be mitigated using structured output formats such as JSON or XML. These formats enable us to enforce consistency and completeness in the model’s output by explicitly defining the expected response structure [12, 22]. In our implementation, we prompt the LLM to respond using JSON schemas, which improves reliability during post-processing and facilitates automatic evaluation of model outputs.
3.1 Experimental Results
Table 1: Comparison PURE with Baselines. We evaluate performance under two data settings: using only item interactions and using item interactions augmented with reviews. $\dagger$ indicates customized baselines where review data is naively incorporated into the original prompt templates designed for item interactions only.
@@ 3.1 Experimental Results
Impact of Review Extractor. Table 1 compares PURE with (1) three baselines solely based on purchased items; (2) modified baselines, marked with $\dagger$, that additionally utilize users’ raw reviews. The results reveal that baselines that simply combine item interactions with raw reviews show inconsistent performance improvements. In contrast, PURE, which leverages the review extractor and profile updater, significantly outperforms all baselines. This demonstrates that processing reviews at three levels, i.e., like, dislike, and key features, is essential for enhancing performance.
Component-wise Study. Table 2 shows the ablation study of PURE, where we analyze the impact of reviews (using or not using) and the effect of components (enabling or disabling the review extractor and profile updater). The use of reviews bring high performance gains only when accompanied by Review Extractor (Ext.). This is due to the sharp increase in input tokens (see the |$T$| column of the 2nd and 3rd rows of each method) as the user continues purchases.
Notably, the best recommendation performance is achieved when Profile Updater (Upd.) is disabled (see the 3rd and 4th rows for each method). This suggests that the well-structured context provided by the Review Extractor alone can lead to strong performance when directly concatenated, even without profile updating. However, it may face a challenge, as the number of purchases grows, leading to significant computational overhead. Thus, we use Profile Updater (Upd.) to maintain compact user profiles, reducing input token size by 15–20% with only a slight 1–3% performance drop. This trade-off underscores the need for Profile Updater for long-term recommendations.
Figure 2: Trade-off between NDCG and input token size.
@@
Trade-off Analysis. We categorize users into three groups based on the total cumulative review token count per user, as the criterion: 0–500 (short), 500–1000 (middle), and 1000–2000 (long) tokens. Figure 2 presents the trade-off between recommendation performance and input token length of the three models including PURE.
PURE achieves the best trade-off, showing the steepest NDCG increase compared to other methods as input token size grows. Therefore, this demonstrates that PURE accurately distills key information from long reviews, while achieving efficiency by minimizing input token growth without information loss, even for long-group users.
4 Related Works
Recommendation Setup. Conventional sequential recommendation methods [23, 10, 24, 25, 26] have followed a one-shot prediction setup, in which a user’s interaction history is split such that the most recent item is held out as the test set, the second-most recent as the validation set, and the remaining history is used for training. While this setup simplifies the evaluation pipeline, it restricts the model to predicting a single target item, thereby failing to capture the nuanced and evolving nature of user preferences over time.
LLM-based Recommendation. A notable example is EXP3RT [27], which constructs static user profiles by fine-tuning LLMs directly on target recommendation datasets. These fine-tuned models are then used to compute preference scores over candidate items. Tallrec [22] proposed the parameter-efficient finetuning (PEFT) method in recommender system, and A-LLMRec [26] proposed to finetune the embedding model for LLM to leverage the collaborative knowledge. In contrast, train-free models [28] guided users through a conversational process to elicit responses and extract multiple features. These features are then used to make personalized recommendations in a conversation-based recommendation framework. Also, uncovering ChatGPT’s capabilities of recommendation  [29] shows ChatGPT is good at reranking the candidates and choosing user preference items while less good at rating. InstructRec [30] designed the instruction to recognize the users’ intention and preference from context.
@@ 5 Conclusion
We propose PURE, a train-free LLM-based recommendation system that operates within a limited input token budget, while maintaining flexibility and eliminating the need for task-specific training. PURE reduces computational costs compared to train-based systems and adapts to various domains. Additionally, we introduce a sequential recommendation task to better model the evolving nature of user preferences over time, moving beyond conventional sequential setups. This work highlights the potential of train-free LLMs in real-world recommendation scenarios.
@@
Limitations. A notable limitation of our approach is the tendency of the LLM to exhibit hallucination by occasionally recommending items beyond the predefined candidate set, even when explicitly instructed to select from it. This phenomenon underscores the difficulty in imposing strict constraints within LLM-based recommendation models while maintaining flexibility and accuracy. Also, our study was constrained by the inability to utilize datasets containing a larger number of user reviews, which may have provided richer context.
Acknowledgements.